Suppose the probability that a library patron checks out a love book is 0.08.

What is the probability that 3 of the next 12 library patrons will sign out a love book? = 0.0532

What is the probability that at least one patron will sign out a love book? = ?????????

The answer is NOT 0.5656, 0.6323, 0.08

What is the probability that none of the next 10 patrons sign out a love book? = 0.4344
Answer #1

Ans:

Use binomial distribution with n=12 and p=0.08

1)P(x=3)=12C3*0.08^3*(1-0.08)^9=0.0532

2)P(at least one)=1-P(none)

=1-P(x=0)

=1-(1-0.08)^12=0.6323

3)

P(none)=P(x=0)=(1-0.08)^10=0.4344